a firm sells a 30-year bond for a discount o 15,000 over its 200,000 face value. if the bond's coupon rate is 7% and they use straight line amortization methods for all intangible assets, what is their interest expense on the bond each year?
Answer #1

Interest expense = Interest paid+Discount amortization

                            = (200000*7%)+(15000/30)

Interest expense = 14500
Answer #2
(200,000x0.07) + (15,000/30) = 14,500
answered by: Bella Sonnek
Answer #3

SOLUTION :


Face value of the bond = 200000 


Discount = 15000 


Maturity period of the bond = 30 years.


Coupon rate is 7% annually.


Hence, coupon amount per year 

= Interest expense on the bond

= 200000 * 0.07

= 14000 


Discount of 15000 is to be amortised for 30 years


As per straight line method amortisation per year

= 15000 / 30

= 500


So, total expense (considered as interest) per year

= Coupon amount + amortised discount 

= 14000 + 500

= 14500 (ANSWER). 

answered by: Tulsiram Garg